# Work statistics for Quantum Spin Chains: characterizing quantum phase transitions, benchmarking time evolution, and examining passivity of quantum states

### v3

@@ Appendix B More results for examining the passivity of ground states in Ising-like and Haldane-like chains
In this Appendix, we present more results of examining the passivity of ground states for Haldance-like chains and for the more non-hermitian actions of Isng-like chains. The first is the plot of $\overline{w}^{\rm impulse}$ in Fig. 15 as a function of coupling parameter $D$ labeling the ground states of the Haldane-like chains under various hermitian actions for small value of $\lambda=0.1$. The second is the plot of $\overline{w}^{\rm impulse}$ as a function of $\lambda\in[-10,10]$ in Fig. 14 to show the passivity of ground states for both hermitian and non-hermitian actions under the cyclic impulse process for the Haldane-like chain, and the third in Fig. 14 is the plot of $\overline{w}^{\rm impulse}$ of Ising-like chains as a function of $\lambda\in[-10,10]$ for the more general non-hermitian actions. The key features are similar to what we have observed in Fig. 9 and Fig. 10 of the main text. In particular, the ground states are all passive for all the cases considered. Moreover, the features shown in Fig. 14 are kind of the mixtures of the top and bottom rows in Fig. 10 as the $V$s considered here are the mixtures of those twos.

### v4

@@ Appendix C More results for examining the passivity of ground states in Ising-like and Haldane-like chains
Then, we present the Haldane-chain counterparts of Fig. 9 and Fig. 10, respectively, in Fig. 16 and Fig. 17. Fig. 16 shows the plot of $\overline{W}^{\rm impulse}$ as a function of coupling parameter $D$ labeling the ground states of the Haldane-like chains under various Hermitian actions for small value of $\lambda=0.1$. Fig. 17 shows the plot of $\overline{W}^{\rm impulse}$ as a function of $\lambda$ to examine the passivity of the ground states under either the Hermitian or non-Hermitian actions of a cyclic impulse process. The key features are similar to what we have observed in Fig. 9 and Fig. 10 of the main text for the Ising chains. In particular, the ground states are all passive for all the cases considered. Moreover, the features shown in Fig. 17 are kind of the mixtures of the top and bottom rows in Fig. 10 as the $V$’s considered here are the mixtures of those twos.
Figure 16: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states of $n=1024$ Haldane-like chain with $J_{z}=1.5$ but varying $D\in[0,2]$ under impluse process of $\lambda=0.1$ and Hermitian actions $V=\sum_{k}S_{k}^{x}$ (black triangles), $\sum_{k}S_{k}^{y}$ (yellow right-arrows), $\sum_{k}S_{k}^{z}$ (red circles), $\sum_{k}(S_{k}^{x})^{2}$ (green diamonds), $\sum_{k}(S_{k}^{y})^{2}$ (orange left-arrows) and $\sum_{k}(S_{k}^{z})^{2}$ (blue squares). The $\overline{W}^{\rm impluse}$ plots show that the ground states are always passive for all considered $V$’s. The degeneracy of $\overline{W}^{\rm impluse}$ under swapping of $S_{i}^{x}$ and $S_{i}^{y}$, as shown here, can be understood as the $Z_{2}$ symmetry of $H_{0}$ and $SO(3)$ spin algebra under the transformation: $(S_{k}^{x},S_{k}^{y},S_{k}^{z})\rightarrow(-S_{k}^{y},S_{k}^{x},S_{k}^{z})$. Besides, $\overline{W}^{\rm impluse}$ always vanishes for $V=\sum_{k}S_{k}^{z}$ because of $[H_{0},\sum_{k}S_{k}^{z}]=0$.
Figure 17: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=1024$ Haldane-like chain with $J_{z}=1.5$. We vary the coupling $D\in[0,2]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of Eq. (11) and Eq. (13) for the Hermitian actions (a) $V=\sum_{k}S_{k}^{x}$, (b) $V=\sum_{k}(S_{k}^{x})^{2}$, (c) $V=\sum_{k}(S_{k}^{z})^{2}$ in the top row, and non-Hermitian actions (d) $V=i\sum_{k}S_{k}^{x}$, (e) $V=\sum_{k}(S_{k}^{x}+iS_{k}^{y})$, and (f) $V=i\sum_{k}(S_{k}^{x})^{2}$ in the bottom row.